OpenStudy (anonymous):

A quiz consists of 10 questions. Each question is marked either “right” or “wrong.” For each student in the class, the instructor keeps a record of the number of right answers the student got. In the class, the average number of right answers is 6.2 and the SD of the number of right answers is 1.4 what will be avg of wrong ans

Directrix (directrix):

@salman_09 Would you read the following and determine what you think about it? Thanks. Insight from @Phone_A_Friend_SS_from_California "I think the answer is 1.4. If x is the number right score than 10-x would be the number wrong score. If M is the st. dev. of right scores then 10-M is the mean of the wrong scores. (assuming here that any non-right response, including leaving answer space blank, is a wrong score. The sd. dev. formula for number right involves summing (x-M)^2. The sd. dev. formula for number wrong scores involves summing [(10-x) - (10-M)]^2 = (-x+M)^2 = (x-M)^2."

OpenStudy (kropot72):

Let the number of students = s Then the number of questions answered = 10s Let the number of right answers = r The average number of questions answered correctly per student = r/s The number of questions answered incorrectly = 10s - r The average number of wrong answers is\[\frac{10s-r}{s}=10-\frac{r}{s}=10-6.2=3.8\]

OpenStudy (anonymous):

"...the average number of right answers is 6.2 . . . what will be avg of wrong ans" That teacher must have been rather sleepy when s/he wrote that question. :-)